OpenStudy (anonymous):

A planet is in a circular orbit around the sun. Which two forces acting on the planet are equal?

OpenStudy (superhelp101):

for sure Gravity and Centrifugal Force. Centrifugal Force is an apparent force that acts outward on a body moving around a center, arising from the body's inertia.

OpenStudy (superhelp101):

got it?

OpenStudy (anonymous):

thanks and do u know what the equation is to demonstrate this relationship?

OpenStudy (anonymous):

@superhelp101

OpenStudy (superhelp101):

equation? honestly no, i am really sorry!

OpenStudy (anonymous):

no worries!

OpenStudy (abhisar):

\(\huge\sf \frac{GMm}{r^2}=\frac{mV^2}{r}\)

OpenStudy (abhisar):

M=mass of sun, m=mass of planet, r=distance between planet and sun, v=velocity of planet.

OpenStudy (vincent-lyon.fr):

Though the expected answer is "Gravity and Centrifugal Force", this question denotes a poor understanding of Newton's laws of motion by those who asked it. If the planet is MOVING about the Sun ie the Sun is fixed, then there is only ONE force acting on the planet : the gravitational force by the Sun on the planet. Saying that there is a centrifugal force acting on the planet IN A CIRCULAR MOTION contradicts Newton's first law of motion => the planet should go in a straight line, not a circle. The centrifugal force only "exists" to explain why the planet is FIXED in a rotating frame of reference comprising the Sun's and the planet's centres.